# The Algorithmic Regulator

### v1

@@ Abstract
The regulator theorem states that, under certain conditions, any optimal controller must embody a model of the system it regulates, grounding the idea that controllers embed, explicitly or implicitly, internal models of the controlled. This principle underpins neuroscience and predictive brain theories like the Free-Energy Principle or Kolmogorov/Algorithmic Agent theory. However, the theorem is only proven in limited settings. Here, we treat the deterministic, closed, coupled world-regulator system ($W$,$R$) as a single self-delimiting program $p$ via a constant-size wrapper that produces the world output string $x$ fed to the regulator. We analyze regulation from the viewpoint of the algorithmic complexity of the output, $K(x)$. Intuitively, knowing that $K(x)$ is small under regulation prunes the class of minimal candidate worlds, but this cannot, in general, certify mutual algorithmic information of world and regulator. This motivates a probabilistic approach. We define $R$ to be a good algorithmic regulator if it reduces the algorithmic complexity of the readout relative to a null (unregulated) baseline $\varnothing$, i.e., $\Delta=K\big(O_{W,\varnothing}\big)-K\big(O_{W,R}\big)>0.$ We then prove that the larger $\Delta$ is, the more world-regulator pairs with high mutual algorithmic information are favored. More precisely, a complexity gap $\Delta>0$ yields $\Pr((W,R)\mid x)\leq C\,2^{\,M(W{:}R)}\,2^{-\Delta}$, making low M(W:R)M(W{:}R) exponentially unlikely as $\Delta$ grows. This is an AIT version of the idea that “the regulator contains a model of the world.” The framework is distribution‑free, applies to individual sequences, and complements the Internal Model Principle. Beyond this necessity claim, the same coding‑theorem calculus singles out a canonical scalar objective and implicates a planner. On the realized episode, a regulator behaves as if it minimized the conditional description length of the readout.
@@ 1 Introduction
The definition ground on mutual algorithmic information $M$ is further motivated by the following: (i) Machine invariance: $M$ is invariant up to $O(1)$ under changes of universal machine. (ii) Distribution‑free: $M$ is defined for individual objects (programs), not probabilistic models. (iii) Operational meaning: M(W:R)M(W{:}R) is precisely the codelength reduction in describing $W$ when $R$ is known, aligning with MDL/Occam reasoning via the Coding Theorem [33, 44, 22]. (iv) Complement to IMP: Unlike the IMP—which requires a dynamical replica of the exosystem under structural assumptions—M(W:R)M(W{:}R) captures any computable shared structure evidenced by contrastive compressibility.
@@
We score regulation by how compressible a task-weighted readout is. Let $x_{t}$ be the weighted error, and let $x_{1:T}$ denote the $T$-sample string. Fix a prefix-free lossless code (e.g., a universal compressor), and define the per-sample code length $L_{T}:=\tfrac{1}{T}\,L_{C}(x_{1:T})$. A regulator $R$ is better on horizon $T$ when it makes $L_{T}$ smaller than a null baseline $\varnothing$, i.e., when the contrastive gap $\Delta:=L_{T}(x;\,\varnothing)-L_{T}(x;\,R)$ is positive. This choice is natural: for stationary ergodic data, the normalized code length of any universal compressor converges to the Shannon entropy rate $h(x)$: one has (almost surely) $\lim_{T\to\infty}K(x_{1:T})/T=h(x)$ [9, 22], so the Kolmogorov‑based criterion reduces to the Shannon criterion whenever those stochastic assumptions are appropriate—while remaining meaningful outside them. For Gaussian processes,
@@
so attenuating the in‑band sensitivity (reducing $S_{xx}$ where it matters [42, 16]) reduces codelength. In short, “compressible error” coincides with the classical intuition that good regulation removes variability/uncertainty in the task band and the IMP [4, 14].

### v2

@@ Abstract
The regulator theorem states that, under certain conditions, any optimal controller must embody a model of the system it regulates, grounding the idea that controllers embed, explicitly or implicitly, internal models of the controlled. This principle underpins neuroscience and predictive brain theories like the Free-Energy Principle or Kolmogorov/Algorithmic Agent theory. However, the theorem is only proven in limited settings. Here, we treat the deterministic, closed, coupled world-regulator system ($W$,$R$) as a single self-delimiting program $p$ via a constant-size wrapper that produces the world output string $x$ fed to the regulator. We analyze regulation from the viewpoint of the algorithmic complexity of the output, $K(x)$ (regulation as compression). We define $R$ to be a good algorithmic regulator if it reduces the algorithmic complexity of the readout relative to a null (unregulated) baseline $\varnothing$, i.e., $\Delta=K\big(O_{W,\varnothing}\big)-K\big(O_{W,R}\big)>0.$ We then prove that the larger $\Delta$ is, the more world-regulator pairs with high mutual algorithmic information are favored. More precisely, a complexity gap $\Delta>0$ yields $\Pr((W,R)\mid x)\leq C\,2^{\,M(W{:}R)}\,2^{-\Delta}$, making low M(W:R)M(W{:}R) exponentially unlikely as $\Delta$ grows. This is an AIT version of the idea that “the regulator contains a model of the world.” The framework is distribution‑free, applies to individual sequences, and complements the Internal Model Principle. Beyond this necessity claim, the same coding‑theorem calculus singles out a canonical scalar objective and implicates a planner. On the realized episode, a regulator behaves as if it minimized the conditional description length of the readout.
@@ 1 Introduction
The definition ground on mutual algorithmic information $M$ is further motivated by the following: (i) Machine invariance: $M$ is invariant up to $O(1)$ under changes of universal machine. (ii) Distribution‑free: $M$ is defined for individual objects (programs), not probabilistic models. (iii) Operational meaning: M(W:R)M(W{:}R) is precisely the codelength reduction in describing $W$ when $R$ is known, aligning with MDL/Occam reasoning via the Coding Theorem [36, 48, 24].
@@
Regulation as compression
We score regulation by how compressible a task‑weighted error stream is. Let $x_{t}$ be the weighted error and $x_{1:T}$ the TT‑sample string. Fix a prefix‑free lossless code (e.g., a universal compressor) and define the per‑sample codelength $L_{T}:=\tfrac{1}{T}L_{C}(x_{1:T})$. A regulator $R$ is better on horizon $T$ when it makes $L_{T}$ smaller than a null baseline $\varnothing$, i.e., when the contrastive gap $\Delta:=L_{T}(x;\varnothing)-L_{T}(x;R)$ is positive. This choice is natural: for stationary ergodic data, normalized universal codelengths converge (a.s.222“Almost surely”: with probability $1$.) to the Shannon entropy rate $h(x)$, and (under standard computability assumptions) $K(x_{1:T})/T=h(x)+o(1)$ almost surely; thus the Kolmogorov‑based criterion reduces to the Shannon criterion when those stochastic assumptions hold—while remaining meaningful outside them [46, 10, 24].
To see the connection between regulation and compression in more detail, let $h_{x_{1:T}}(\alpha):=\min_{S\ni x_{1:T},\,K(S)\leq\alpha}\log|S|$ denote the Kolmogorov structure function [41]. Regulation amounts to moving down this curve: as the regulator invests model bits (larger $\alpha$), more regularity in $x_{1:T}$ is captured and the residual randomness $h_{x_{1:T}}(\alpha)$ drops, approaching 0 at perfect regulation. The notion of robability emerges along this path. Replacing set‑models by probabilistic models $\{P_{M}\}$ turns the two‑part description into the standard MDL form
$L(x_{1:T};M)\;\approx\;K(M)\;-\!\sum_{t=1}^{T}\log P_{M}(x_{t}),$
where the second term is the ideal codelength under $P_{M}$ (Shannon coding: $\log P_{M}$) [10, 24, 18]. If the regulator must hedge over multiple models, the mixture/Bayes code with prior $\pi$ uses $\bar{P}(x_{1:T})=\sum_{M}\pi(M)P_{M}(x_{1:T})$ and assigns
$L_{\rm mix}(x_{1:T})\;=\;-\log\bar{P}(x_{1:T})\;=\;-\log\!\sum_{M}\pi(M)P_{M}(x_{1:T}),$
a valid prefix code whose regret relative to the best single model $M^{\star}$ is bounded by $\log\pi(M^{\star})$; with $\pi(M)\!\propto\!2^{-K(M)}$ (Solomonoff/Occam), the penalty matches the model description length [6, 18, 35, 36, 24]. Thus, probabilistic/Bayesian regulation is the coding‑optimal way to descend $h_{x_{1:T}}(\alpha)$, aligning with the multi‑model argument in [29].
Finally, we can treat the regulator input as an error signal quantized at fixed sensor resolution; the per‑sample codelength of $x_{1:T}$ under a universal compressor converges to the entropy rate for stationary sources. For Gaussian processes,
@@
where $S_{xx}$ is the input power spectral density [46, 17] of the error signal $x$, so attenuating in‑band sensitivity (reducing $S_{xx}$ where it matters) reduces codelength [17]. In the scalar white‑Gaussian case with variance $\sigma^{2}$, $h=\tfrac{1}{2}\log(2\pi e\,\sigma^{2})$, so smaller‑amplitude fluctuations (smaller $\sigma$) mean lower entropy and better compressibility. In short, “compressible error” matches the classical view: good regulation removes variability/uncertainty in the task band and accords with the IMP [4, 15].